# Data Management in the Noisy Intermediate-Scale Quantum Era

### v1

@@ 1. Introduction
  • Fundamentals: We explain quantum data, differentiate it from classical data, and describe the transformation processes between them (Sec. 2).
@@
  • Research problems: We elaborate on near-term research questions in data management for quantum computing (Sec. 4).
2. Data in quantum era
Since the 1960s, database management systems (DBMSs) have been developed and operated on classical computers, which rely on bits as fundamental units of data. Classical bits represent binary values of 0 or 1. In the development of quantum computers and their associated software systems, a key challenge is the implementation of qubits (Popkin, 2016), the quantum counterparts of classical bits. Following the principles of quantum mechanics, qubits differ from bits in properties such as superposition and entanglement. These properties fundamentally differentiate quantum data from classical data, which we will explore in this section.
2.1. Classical data vs. quantum data
@@ 2. Data in quantum era
1. Quantum data is probabilistic. Unlike a classical bit, which is 0 or 1, a quantum bit can be in superposition. Mathematically, a zero state and a one state may be represented by a unit vector in the standard basis. That is, the zero state |$\ket{0}$ is represented by the vector $\begin{bmatrix}1\\ 0\end{bmatrix}$ and the one state |$\ket{1}$ is represented by the vector $\begin{bmatrix}0\\ 1\end{bmatrix}$.  A single qubit state, denoted |$\ket{\psi}$, may be represented by a superposition, i.e., a linear combination of |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, for some pair of complex numbers $\alpha,\beta$ called amplitudes, which satisfy $|\alpha|^{2}+|\beta|^{2}=1$.  The probabilistic nature arises when a measurement is performed. When a measurement is performed on the state |$\ket{\psi}$, an outcome 0 is obtained with probability $|\alpha|^{2}$ and 1 obtained with probability $|\beta|^{2}$, and the state permanently changes to the obtained outcome.
Example 2.1 (Superposition).
Consider the qubit in the following superposition:
|$\vspace{-0.2cm}\ket{\psi}=\frac{1}{\sqrt{2}}\ket{0}+\frac{1}{\sqrt{2}}\ket{1}$
When measuring this qubit, there is an equal probability of 50% ($\left|\alpha\right|^{2}=\left|\beta\right|^{2}=\frac{1}{2}$) to get a 0 or 1 as a result.
In classical computing, individual bits can be concatenated to form a bit string, e.g., the three-bit string “010.” Similarly, we may concatenate qubits into a multi-qubit state. In this case, an $n$-qubit state can be represented as a superposition of |$\ket{x}$ for $x\in\{0,1\}^{n}$, or equivalently a vector of $2^{n}$ components.  Later in Sec. 4, we will discuss a three-qubit state in Fig. 2a.
@@
3. Quantum data can be entangled. Another difference between qubits and bits is entanglement (Einstein et al., 1935), leading to spooky action at a distance, a phrase coined by Einstein. Imagine two qubits with qubit $A$ being in Amsterdam and qubit $B$ in San Francisco. When qubit $A$ and qubit $B$ are entangled this means their states are correlated. The spooky part is that if we measure qubit $A$ in Amsterdam and find it in a state of 00 (or $1$), we instantaneously know the state of qubit $B$ is 00 (or $1$), even though it is far away in San Francisco. This correlation is maintained independently of the physical distance between them. The next example introduces Bell’s state, a well-known entangled state.
Example 2.2.
(Bell’s state) We extend the one-qubit state in Example 2.1 to a two-qubit state as follows:
|$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})=\frac{1}{\sqrt{2}}\begin{bmatrix}1\\ 0\\ 0\\ 1\end{bmatrix}\vspace{-0.1cm}$
2.2. Quantum-classical data transformation
The transformations that map quantum data to classical data, and vice versa, are essential ingredients of quantum computing.  These transformations can be described as quantum algorithms.
Quantum algorithms, circuits, and gates (Nielsen and Chuang, 2010). A quantum algorithm is a carefully designed quantum process that maps input data, classical or quantum, to the desired output data.  The most commonly used model to describe quantum algorithms is the quantum circuit model. A quantum circuit is a sequence of operations, quantum gates, which manipulate qubits to perform computations. Quantum gates are unitary operators. An operator $U$ is unitary if its conjugate transpose $U^{\dagger}$ multiplied by $U$ (and vice versa) equals the identity matrix $I$, denoted as $U^{\dagger}U=UU^{\dagger}=I$. A unitary operator may be composed as a sequence of basic unitary operators. A set of quantum gates is universal if it can approximate any unitary operator to desirable precision.  A widely known universal gate set consists of Pauli gates (including $X,Y,Z$ gates defined by Pauli matrices), the Hadamard gates (see Example 2.3), the $T$ gate, and the controlled-NOT (CNOT) gate.
Example 2.3.
The Hadamard gate (H) is a unitary operator, which maps |$\ket{0}$ to (|$(\ket{0}+\ket{1})/\sqrt{2}$, and |$\ket{1}$ to (|$(\ket{0}-\ket{1})/\sqrt{2}$.  In matrix form, the operator is represented as
$H=\frac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]$
2.2.1. From quantum data to classical data
To solve computational problems using quantum computers, it is essential to extract classical information from quantum states.  In quantum physics, this is done through measurements, which manipulate and convert quantum states (quantum data) into classical outcomes (classical data). For example, measuring the quantum state |$\ket{\psi}$ in Example 2.1 yields a result of 00 with probability $|\alpha|^{2}$ and $1$ with probability $|\beta|^{2}$.
2.2.2. From classical data to quantum data
To load classical data into a quantum computer, the necessary step is encoding the classical data into the states of qubits. Classical data can be encoded as quantum states via encoding methods like basis encoding, amplitude encoding, and many advanced methods (Marin-Sanchez et al., 2023; Weigold et al., 2021a, b, 2022; Nakaji et al., 2022). Basis encoding maps classical bits of 0 and 1 directly onto the quantum states |$\ket{0}$ and |$\ket{1}$. Amplitude encoding encodes classical data into the amplitudes of quantum states. For instance, we have classical data, a normalized vector $x=\begin{bmatrix}a\\ b\end{bmatrix}=\begin{bmatrix}\frac{3}{5}\\ \frac{4}{5}\end{bmatrix}$. We can encode $x$ as a quantum state $|\psi\rangle=a|0\rangle+b|1\rangle=\frac{3}{5}|0\rangle+\frac{4}{5}|1\rangle$.
@@ 3. Our Vision: Data Management for Quantum Computing
To facilitate future research on data management for quantum computing, we first sketch the whole landscape in Fig. 1. We categorize this landscape into three distinct paradigms based on how quantum and classical data are transformed and utilized, and the type of hardware involved—whether a quantum or classical computer is employed. Each paradigm not only guides the exploration of quantum research topics leveraging database technologies but also paves the way for new opportunities in database systems, ultimately driving the development of quantum computing.
I Classical paradigm: Classical preprocessing & postprocessing for quantum computing
@@
We begin with the most prevalent setting today: classical computers handle the preprocessing and postprocessing of data for quantum devices, such as quantum chips, quantum sensors, and quantum computers. In this paradigm, data is primarily stored and processed as classical data. For instance, a classical computer may send instructions to a quantum chip. These instructions are classical data and, for example, describe the quantum circuit that should be run. The quantum device runs its computation and, after some time, performs measurements onto one or more of its qubits. The measurement outcomes, which are bits and hence classical data, are then returned to the classical computer. Thus, the quantum device receives classical input (the instructions) and returns the classical output (the measurement outcomes). For some applications, this process is continued as a loop: the measurement output of the current iteration contributes to the next iteration’s input instructions, and both are classical data.
@@
Database research can enhance this process by developing efficient systems for managing, storing, and querying the classical data involved in preprocessing, postprocessing, and iterative feedback loops between classical computers and quantum devices, ensuring seamless integration and optimization of data workflows.
@@
We demonstrate the importance of this paradigm through two key categories of quantum applications: first, applications where the quantum computation is completed immediately upon returning a result. For example, two separated quantum computers can generate a secure, shared key (password) by encoding bits into qubits, then sending and measuring the qubits  (Bennett and Brassard, 1984; Pirandola et al., 2020). This is followed by classical postprocessing on the bitstring (i.e. classical data) that the measurement returns (Elkouss et al., 2011). The resulting bitstring is a secret key that can later be used for secure communication. In the other category, the quantum computation is stalled temporarily, and the quantum chip still holds quantum data on which the computation will continue later. An example is the detection of errors during a quantum computation, where at fixed timesteps during the computation check measurements are performed (Gottesman, 1997; Terhal, 2015), whose outcomes are then decoded to find out which error has most likely occurred (Maurya and Tannu, 2024; Battistel et al., 2023).
@@
In this paradigm, often, fast processing of classical data (including I/O with quantum devices) is crucial for practical use, as quantum devices in the NISQ era are noisy, and the quality of quantum data degrades rapidly over time. We delve into specific research challenges, such as error correction using database technologies, in Sec. 4.3.
II Classical-quantum paradigm: classical data represents quantum states and operations.
This hybrid paradigm presents significant potential for new database challenges. It focuses on using classical data to represent and simulate quantum data. For example, Bell’s state from Example 2.2 is represented by a classical vector. Research in this paradigm is more accessible because it does not require a quantum computer; instead, it involves classical computers.
The representative task in this hybrid paradigm is simulation. Simulation is the process of emulating quantum computation, enabling researchers to model and analyze quantum processes as if they were operating on actual quantum hardware (Young et al., 2023; Xu et al., 2023). [^1] In this work, by simulation we refer to classical simulation. Another related term is quantum simulation (Buluta and Nori, 2009; Georgescu et al., 2014), which pertains to simulating quantum mechanics on a quantum computer, a subject studied in physics. Simulations are of paramount importance in the NISQ era (Preskill, 2018). Consider, for instance, the concept of quantum supremacy (Boixo et al., 2018), which seeks to demonstrate the superior capabilities of quantum computers over classical computers. Given that large-scale quantum computers are not yet available in the NISQ era, simulation is essential for comparing the scalability of quantum computers with classical ones. Additionally, simulations play a crucial role in the development of new quantum algorithms, allowing researchers to design, debug, and validate the correctness of these algorithms before deploying them on expensive quantum devices. Moreover, simulations aid in the development of quantum hardware by evaluating error mitigation schemes, predicting algorithm runtimes, and more (Zulehner and Wille, 2018; Jones et al., 2019; Zhou et al., 2020b; Villalonga et al., 2019; Azuma et al., 2021). Simulation serves as a foundational tool across key areas of quantum computing, including quantum supremacy, quantum algorithms, quantum hardware, error correction, and the exploration of potential quantum applications (Xu et al., 2023).
The potential for database research in this paradigm is immense. First, imagine innovative database systems specifically designed to manage classical data that supports simulation. Such systems could handle complex queries, such as updating quantum-state representations as gates are applied, returning classical measurement outcomes, or calculating the closeness of the current quantum state to a desired target state (Vinkhuijzen et al., 2024).
Second, efficient caching strategies for expensive operations, such as repeated simulations of quantum circuits with varying error parameters, present another important research direction. These strategies require the development of efficient data management systems that optimize the computation of such operations.
The third research direction in this paradigm involves the representation of quantum states and operations. Not only can a quantum state be represented as classical data (e.g., vectors), but quantum gates can also be represented as classical data, such as matrices  (van de Wetering, 2020; Hubig et al., 2017). This is useful for quantum design questions such as optimizing the number of quantum gates or gate depth (Kissinger and van de Wetering, 2020; Zulehner et al., 2018) and compiling textbook quantum circuits to only include the low-level operations that real quantum devices support (Smith et al., 2020). As the representation of a quantum state (a vector) or quantum gate (a matrix) generally grows exponentially with the number of qubits, developing efficient methods to store and process these data structures is key to making simulations of quantum computation feasible, either in vector form or through more compressed structures like tensor networks (Orús, 2019) or other advanced representations (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
As we delve deeper into this paradigm, we explore specific challenges and database research opportunities in Sec. 4.1 and 4.2, where we outline the paths forward for leveraging database techniques to meet the emerging demands of quantum computing.
@@
III Quantum paradigm: quantum-native data storage and processing.
@@
Finally, we envision a future research paradigm beyond NISQ, when we have large-scale, fault-tolerant quantum computers. We will deal with pure quantum data, i.e., qubits. Here, the quantum hardware supports storing qubits for long enough to perform other tasks in the meantime. This enables for example a cloud quantum computer (Broadbent et al., 2009) who rotates its resources among multiple clients, various types of quantum machine learning (Cerezo et al., 2022), and Quantum Random Access Memory (Giovannetti et al., 2008), where data can be queried in superposition. A recent vision (Huang et al., 2022) shows a possible future: quantum data is collected from quantum sensing systems, e.g., for discovering a black hole; then stored and processed via the quantum memory of a quantum computer.
@@ 4. Near-Term Research Problems
We will now explore data management research questions across the three paradigms. In Sec. 4.1 and 4.2 we focus on a key challenge within the classical-quantum paradigm. In Sec. 4.3 we broaden the scope to include research opportunities across all three paradigms.
Figure 2. (a) The 3-qubit GHZ state $\frac{1}{\sqrt{2}}(\ket{000}+\ket{111})$ (Nielsen and Chuang, 2010) represented in state vector, whose size is $82^{3}=$ 8. (b) GHZ state represented as a tensor network (Pérez-Garcia et al., 2007; Orús, 2014). For better visualization, in the grey box, we write the tensors in vector/matrix form by slicing the last dimension of A, B, and C. (c) Tensor contraction between matrices $D(i\times k)$ and $E(k\times j)$, leading to a new tensor $F(i\times j)$. (d) The Hadamard gate H (Example 2.3) represented in tensor network. During simulation, the gate operation $H$ executed on qubit state $G$ is represented by contracting $G$ and $H$, leading to a new state $G^{\prime}$.
@@ 4.1. The challenge of simulating quantum computation
The input for a simulation is a quantum algorithm described as a quantum circuit, [^2] A uniform family of circuits, to be precise, one for each input size. consisting of input qubits upon which quantum gates are applied, ultimately resulting in a probability vector of the output states, i.e., measurement outcomes (see Sec.  2.2). Then, a simulation is a classical computation that computes that output distribution (so-called strong simulation), or samples from it (weak simulation) (Van Den Nes, 2010). We primarily focus on strong simulation.

### v2

# Quantum Data Management in the NISQ Era: Extended Version

@@ Abstract.
PVLDB Reference Format:
PVLDB, 18(6): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 6 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/InfiniData-Lab/Quantum.
@@ 1. Introduction
  • Fundamentals: We explain quantum data, differentiate it from classical data (Sec. 2).
@@
  • Research problems: We elaborate on near-term research questions in data management for quantum computing, and report on preliminary experimental results (Sec. 4).
2. Data in the quantum era
@@
1. Quantum data is probabilistic. Unlike a classical bit, which is 0 or 1, a quantum bit can be in superposition. Mathematically, a zero state and a one state may be represented by a unit vector in the standard basis. That is, the zero state |$\ket{0}$ is represented by the vector $\begin{bmatrix}1\\ 0\end{bmatrix}$ and the one state |$\ket{1}$ is represented by the vector $\begin{bmatrix}0\\ 1\end{bmatrix}$.  A single qubit state, denoted |$\ket{\psi}$, may be represented by a superposition, i.e., a linear combination of |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, for some pair of complex numbers $\alpha,\beta$ called amplitudes, which satisfy $|\alpha|^{2}+|\beta|^{2}=1$.  The probabilistic nature arises when a measurement is performed. When a measurement is performed on the state |$\ket{\psi}$, an outcome 0 is obtained with probability $|\alpha|^{2}$ and 1 obtained with probability $|\beta|^{2}$, and the state permanently changes to the obtained outcome. In classical computing, individual bits can be concatenated to form a bit string, e.g., the three-bit string “010.” Similarly, we may concatenate qubits into a multi-qubit state. In this case, an $n$-qubit state can be represented as a superposition of |$\ket{x}$ for $x\in\{0,1\}^{n}$, or equivalently a vector of $2^{n}$ components.
@@
3. Quantum data can be entangled. Another difference between qubits and bits is entanglement (Einstein et al., 1935), which means that multiple qubits are correlated such that measuring the state of one qubit immediately affects the others. A well-known entangled state is the Bell’s state:
|$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1\\ 0\\ 0\\ 1\end{smallmatrix}\right],$
where knowledge of one qubit determines the other. See Appendix A of our online report (Hai et al., 2025) for more background.
@@ 3. Our Vision: Data Management for Quantum Computing
To facilitate future research on data management for quantum computing, we first sketch the whole landscape in Fig. 1. We categorize this landscape into three distinct paradigms based on how quantum and classical data are transformed and utilized, and the type of hardware involved—whether a quantum or classical computer is employed. Our goal is to introduce data management researchers to quantum use cases, and by making a distinction based on the nature and role of data within cases.
I Classical simulation of quantum computing paradigm: classical data represents quantum states and operations.
This paradigm presents great potential for new database challenges. It focuses on using classical data to represent and simulate quantum data. This paradigm is more accessible as it relies on classical computers, not quantum ones.
The representative task in this paradigm is simulation. Simulation is the process of emulating quantum computation, enabling researchers to model and analyze quantum processes as if they were operating on actual quantum hardware (Young et al., 2023; Xu et al., 2023). [^1] In this work, by simulation we refer to classical simulation. Another related term is quantum simulation (Buluta and Nori, 2009; Georgescu et al., 2014), which pertains to simulating quantum mechanics on a quantum computer, a subject studied in physics. Simulations are of paramount importance in the NISQ era (Preskill, 2018). Consider, for instance, the concept of quantum supremacy (Boixo et al., 2018), which seeks to demonstrate the superior capabilities of quantum computers over classical computers. Given that large-scale quantum computers are not yet available in the NISQ era, simulation is essential for comparing the scalability of quantum computers with classical ones. Additionally, simulations play a crucial role in the development of new quantum algorithms, allowing researchers to design, debug, and validate the correctness of these algorithms before deploying them on expensive quantum devices. Moreover, simulations aid in the development of quantum hardware by evaluating error mitigation schemes, predicting algorithm runtimes, and more (Zulehner and Wille, 2018; Jones et al., 2019; Zhou et al., 2020b; Villalonga et al., 2019; Azuma et al., 2021). Simulation serves as a foundational tool across key areas of quantum computing, including quantum supremacy, quantum algorithms, quantum hardware, error correction, and the exploration of potential quantum applications (Xu et al., 2023).
The potential for database research in this paradigm is immense. First, imagine innovative database systems specifically designed to manage classical data that supports simulation. Second, efficient caching strategies for expensive operations, such as repeated simulations of quantum circuits with varying error parameters, present another important research direction. The third research direction in this paradigm involves the representation of quantum states and operations. This is useful for quantum design questions such as optimizing the number of quantum gates or gate depth (Kissinger and van de Wetering, 2020; Zulehner et al., 2018) and compiling textbook quantum circuits to only include the low-level operations that real quantum devices support (Smith et al., 2020). As the representation of a quantum state (a vector) or quantum gate (a matrix) generally grows exponentially with the number of qubits, developing efficient methods to store and process these data structures is key to making simulations of quantum computation feasible, either in vector form or through more compressed structures like tensor networks (Biamonte, 2020; Orús, 2019) or other advanced representations (Bravyi and Gosset, 2016; van de Wetering, 2020a; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
II Joint Quantum-Classical Computing paradigm: classical preprocessing & postprocessing for quantum computing
@@
This paradigm focuses on the situation involved in most quantum-technology applications: classical computers handle the preprocessing and postprocessing of data for quantum devices, such as quantum chips, quantum sensors, and quantum computers. In this paradigm, the focus is on data which is primarily stored and processed as classical data. For instance, a classical computer may send instructions to a quantum chip. These instructions are classical data and, for example, describe the quantum circuit that should be run. The quantum device runs its computation and, after some time, performs measurements onto one or more of its qubits. The measurement outcomes, which are bits and hence classical data, are then returned to the classical computer. Thus, the quantum device receives classical input (the instructions) and returns the classical output (the measurement outcomes).
@@
Database research can enhance this process by developing efficient systems for managing, storing, and querying the classical data involved in preprocessing, postprocessing, and iterative feedback loops between classical computers and quantum devices, ensuring seamless integration and optimization of data workflows. For instance, the development of quantum error correction—a critical area in quantum computing—can benefit significantly from efficient graph data analytics techniques, as further discussed in Sec. 4.3.
@@
We demonstrate the importance of this paradigm through three key categories of quantum applications: first, applications where the quantum computation is completed immediately upon returning a result. For example, two separated quantum computers can generate a secure, shared key (password) by encoding bits into qubits, then sending and measuring the qubits  (Bennett and Brassard, 1984; Pirandola et al., 2020). This is followed by classical postprocessing on the bitstring (i.e. classical data) that the measurement returns (Elkouss et al., 2011). The resulting bitstring is a secret key that can later be used for secure communication. In the second category, the quantum computation is stalled temporarily, and the quantum chip still holds quantum data on which the computation will continue later. An example is the detection of errors during a quantum computation, where at fixed timesteps during the computation check measurements are performed (Gottesman, 1997; Terhal, 2015), whose outcomes are then decoded to find out which error has most likely occurred (Maurya and Tannu, 2024; Battistel et al., 2023). Third, efficient preprocessing is critical for quantum algorithms, which requires loading or encoding classical data into the quantum domain, typically by mapping classical bits onto quantum bits. This process often requires encoding classical data into a quantum circuit, either as part of the algorithm (Sünderhauf et al., 2024) or to output a quantum state with amplitudes representing the classical data (Cortese and Braje, 2018), which is a challenging problem.
@@
III Pure quantum computing paradigm: quantum-native data storage and processing.
@@
Finally, we envision a future research paradigm beyond NISQ, when we have large-scale, fault-tolerant quantum computers. We will deal with pure quantum data, i.e., qubits. Here, the quantum hardware supports storing qubits for long enough to perform other tasks in the meantime. This enables for example a cloud quantum computer (Broadbent et al., 2009) which rotates its resources among multiple clients, various types of quantum machine learning (Cerezo et al., 2022), and Quantum Random Access Memory (Giovannetti et al., 2008), where data can be queried in superposition. A recent vision (Huang et al., 2022) shows a possible future: quantum data is collected from quantum sensing systems, e.g., for discovering a black hole; then stored and processed via the quantum memory of a quantum computer.
@@ 4. Near-Term Research Problems
We will now explore data management research questions across the three paradigms. In Sec. 4.1 and 4.2, we focus on a key challenge within Paradigm I: Classical simulation of quantum computing paradigm. In Sec. 4.3, we broaden the scope to include research opportunities across all three paradigms.
@@ 4.1. The challenge of simulating quantum computation
The input for a simulation is a quantum algorithm described as a quantum circuit, [^2] A uniform family of circuits, to be precise, one for each input size. consisting of input qubits upon which quantum gates are applied, ultimately resulting in a probability vector of the output states, i.e., measurement outcomes (see Sec. 2). Then, a simulation is a classical computation that computes that output distribution (so-called strong simulation), or samples from it (weak simulation) (Van Den Nes, 2010). We primarily focus on strong simulation.
Figure 2. (a) The 3-qubit GHZ state $\frac{1}{\sqrt{2}}(\ket{000}+\ket{111})$ represented as a state vector of size $82^{3}=$ 8. (b) GHZ state as a tensor network. For better visualization, in the grey box, we write the tensors in vector/matrix form by slicing the last dimension of A, B, and C. See Appendices A and B1 of (Hai et al., 2025) for details.
[…]